# Angular integrals with three denominators via IBP, mass reduction, dimensional shift, and differential equations

### v1

@@ III Mass reduction via two-point splitting
$\tilde{I}_{111}^{(3)}$ $(v_{12},v_{13},v_{23},v_{11},v_{22},v_{33})=\tilde{I}_{111}^{(1)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)},v_{33}\right)-\tilde{I}_{111}^{(0)}\left(v_{(12)(13)},v_{(12)(31)},v_{(13)(31)}\right)$
$+\tilde{I}_{111}^{(1)}\left(v_{1(12)},v_{1(13)},v_{(12)(13)},v_{11}\right)+\tilde{I}_{111}^{(1)}\left(v_{3(12)},v_{3(31)},v_{(12)(31)},v_{33}\right)-\tilde{I}_{111}^{(0)}\left(v_{(21)(23)},v_{(21)(32)},v_{(23)(32)}\right)$
$+\tilde{I}_{111}^{(1)}\left(v_{2(21)},v_{2(23)},v_{(21)(23)},v_{22}\right)+\tilde{I}_{111}^{(1)}\left(v_{3(21)},v_{3(32)},v_{(21)(32)},v_{33}\right).$ (22)
IV Dimensional-shift identity
@@
Also the master integrals with $j_{i}=1$ and $v_{ii}=0$, which are collinearly divergent in $d=4$ dimensions, converge for sufficiently large dimension, i.e. small $\varepsilon$. In particular they are finite in $d=6$ dimensions. Hence applying eq.(23) to $I_{111}(\varepsilon)$ and subsequently reducing the integrals in six dimensions on the right to master integrals, results in all $1/\varepsilon$ poles being captured in prefactors of the master integrals.
As it turns out the identity connecting $I_{111}(\varepsilon)$ and $I_{111}(\varepsilon-1)$ simplifies further by shifting all integrals with two or less denominators back to $d=4-2\varepsilon$ dimensions. The necessary identities follow from eq.(23) by subsequently choosing zero, one, and two denominators, performing an IBP reduction and solving for the $d=6-2\varepsilon$ dimensional integrals. Going through the outlined steps, we find the remarkably simple relation
$I_{111}(\varepsilon)=\frac{1}{X_{123}}\left[X_{12\bar{3}}I_{110}(\varepsilon)+X_{1\bar{2}3}\,I_{101}(\varepsilon)+X_{\bar{1}23}\,I_{011}(\varepsilon)+\frac{2\varepsilon}{1-2\varepsilon}\,Y_{123}\,I_{111}(\varepsilon-1)\right].$ (24)
@@
Using the formula (24) has two benefits. Firstly for the purpose of analytical calculation, we gain the pole part and the $\varepsilon^{0}$ term “for free” since they are contained entirely in terms of known two-denominator integrals. Since $I_{111}(\varepsilon-1)$ is finite for $\varepsilon=0$ and is multiplied by an explicit $\varepsilon$ in eq. (24), it only starts contributing at order $\varepsilon$. Hence, for results to order $\varepsilon$ it suffices to calculate the leading term of $I_{111}(\varepsilon-1)$ by differential equations.
Secondly, from a numerical perspective, by expanding the integration measure of $I_{111}(\varepsilon-1)$ in a series in $\varepsilon$ we obtain a finite integral representation for every order in $\varepsilon$. In conjunction with the known expansion of the two-denominator integrals this can be used for numeric checks.
V Differential operators for angular integrals
@@
Derivatives of the other invariants $v_{22}$, $v_{23}$, $v_{33}$ are trivially 00 due to the chosen ansatz. For $\partial/\partial v_{12}$, in the massless case, where $v_{11}=v_{22}=v_{33}=0$, we find
$\frac{\partial}{\partial v_{12}}=\frac{1}{Y_{123}^{(0)}}\Big{[}(v_{13}+v_{23}-v_{12}-v_{13}v_{23})\vec{v}_{1}-(2-v_{13})v_{13}\vec{v}_{2}+(v_{12}+v_{13}-v_{23}-v_{12}v_{13})\vec{v}_{3}\Big{]}\cdot\frac{\partial}{\partial\vec{v}_{1}},$ (38)
@@
$\frac{\partial}{\partial v_{11}}=\frac{1}{2Y_{123}^{(1)}}\Big{[}-v_{23}(2-v_{23})\vec{v}_{1}+(v_{13}+v_{23}-v_{12}-v_{13}v_{23})\vec{v}_{2}+(v_{12}+v_{23}-v_{13}-v_{12}v_{13})\vec{v}_{3}\Big{]}\cdot\frac{\partial}{\partial\vec{v}_{1}},$ (40)
@@
VI Differential equations
Having constructed suitable differential operators we are now equipped to derive differential equations for the three-denominator angular integral. The only master integral with genuinely three denominators is $I_{111}$. All other master integrals have at most two denominators and are known to all orders in $\varepsilon$. To construct the full solution for an arbitrary number of masses, we have to look at two cases. First we consider the massless integral. It is symmetric in its variables $v_{12},v_{13},v_{23}$, hence it suffices to only consider the differential equation $\partial/\partial v_{12}$ and construct a symmetric solution. As a boundary value we will use the integral at the symmetric point $v_{12}=v_{13}=v_{23}=1$. The second case to consider is the single-massive integral with $v_{11}\neq 0$. Here, we look at the differential equation with respect to $\partial/\partial v_{11}$ and use the massless integral as a boundary condition. As discussed in section IV it is beneficial to consider the differential equations for integrals in $d=6-2\varepsilon$ dimensions. This way, the leading-order solution in $\varepsilon$ of the differential equation, which is of order $\varepsilon^{0}$, will provide a contribution to the order $\varepsilon$ of the master integral in $d=4-2\varepsilon$ dimensions.
@@ VI Differential equations
$\frac{\partial I^{(0)}_{111}}{\partial v_{12}}=\frac{1}{Y_{123}^{(0)}}$ $\Big{[}v_{13}(v_{13}-v_{12}-v_{23})I^{(0)}_{211}+(v_{12}-v_{13}-v_{23}+v_{13}v_{23})I^{(0)}_{111}+v_{13}(2-v_{13})I^{(0)}_{201}$
$+(v_{23}-v_{12}-v_{13}+v_{12}v_{13})I^{(0)}_{210}\Big{]}$ (42)
@@
$\frac{\partial I^{(1)}_{111}}{\partial v_{11}}=\frac{1}{2Y_{123}^{(1)}}$ $\Big{[}v_{23}(v_{23}-v_{12}-v_{13})I^{(1)}_{211}+v_{23}(2-v_{23})I^{(1)}_{111}+(v_{13}-v_{12}-v_{23}+v_{12}v_{23})I^{(1)}_{210}$
$+(v_{12}-v_{13}-v_{23}+v_{13}v_{23})I^{(1)}_{201}\Big{]}.$ (43)
From here, we can use the recursion relations from appendix C to reduce the right-hand side of both equations to master integrals. In both cases, this results in a closed differential equation for $I_{111}$, where all other terms are expressed in terms of known lower-denominator integrals.
@@
$\frac{\partial I^{(0)}_{111}}{\partial v_{12}}=\left[\frac{d-6}{2v_{12}}-\frac{d-5}{Y_{123}^{(0)}}\,(v_{13}+v_{23}-v_{23}-v_{13}v_{23})\right]$ $I^{(0)}_{111}+\frac{1}{2v_{12}Y_{123}^{(0)}}\Big{[}(d-5)v_{12}(4-v_{12}-v_{13}-v_{23})I^{(0)}_{110}$
$+(d-5)(2-v_{13})(v_{23}-v_{12}-v_{13})$ $I^{(0)}_{101}+(d-5)(2-v_{23})(v_{13}-v_{12}-v_{23})I^{(0)}_{011}\Big{]}.$ (44)
@@
$\frac{\partial I^{(1)}_{111}}{\partial v_{11}}$ $=\left[\frac{(6-d)v_{23}}{2(2v_{12}v_{13}-v_{11}v_{23})}+\frac{(d-5)v_{23}(2-v_{23})}{2Y_{123}^{(1)}}\right]I^{(1)}_{111}$
$+\frac{d-5}{2Y_{123}^{(1)}(2v_{12}v_{13}-v_{11}v_{23})}\Big{[}\Big{(}(v_{12}+v_{13}-v_{23})(v_{12}+v_{23}-v_{13}-v_{12}v_{23})+Y_{123}^{(1)}\Big{)}I_{110}^{(1)}$
$+\Big{(}(v_{12}+v_{13}-v_{23})(v_{13}+v_{23}-v_{12}-v_{13}v_{23})+Y_{123}^{(1)}\Big{)}I_{101}^{(1)}-(v_{12}+v_{13}-v_{23})v_{23}(2-v_{23})I_{011}^{(0)}\Big{]}$
$+\frac{(d-4)(1-v_{11})}{v_{11}(2v_{12}v_{13}-v_{11}v_{23})}I_{100}^{(1)}+\frac{3-d}{v_{11}(2v_{12}v_{13}-v_{11}v_{23})}I_{000}^{(0)}\,.$ (45)
We note that in both eqs. (44) and (45) all master integrals except for $I_{111}^{(0,1)}$ are known analytically in $d$ dimensions.
As discussed in section IV, it is most useful to consider the differential equations in $d=6-2\varepsilon$ dimensions. The standard procedure to approach differential equations of this form is bringing them to canonical $\varepsilon$-form, i.e. making the coefficient in front of $I_{111}$ proportional to $\varepsilon$. In the particular case at hand we can even make the entire coefficient disappear, leaving us with a right-hand side which is expressed solely in terms of known integrals This is achieved by rescaling the integral to $J_{111}=f(v)I_{111}$. Schematically writing the original eqs. (44) and (45) as
$\frac{\partial I_{111}}{\partial v}\,=\,c(v)\,I_{111}\,+\,r(v)\,,$ (46)
@@
$\frac{\partial J_{111}}{\partial v}\,=\left(\frac{1}{f(v)}\frac{\partial f(v)}{\partial v}\,+\,c(v)\right)J_{111}\,+\,f(v)\,r(v)\,.$ (47)
@@
$J_{111}^{(0)}$ $\equiv(v_{12}v_{13}v_{23})^{\varepsilon}\left(Y^{(0)}_{123}\right)^{1/2-\varepsilon}I_{111}^{(0)}.$ (48)
The prefactor is chosen such that in the differential equation for $\partial/\partial v_{12}J_{111}^{(0)}$ the coefficient in front of $J_{111}^{(0)}$ becomes zero while simultaneously preserving the symmetry between variables. For the single-massive integral we rescale as
$J_{111}^{(1)}$ $\equiv\left(\frac{v_{23}}{2}\right)^{\varepsilon}(2v_{12}v_{13}-v_{11}v_{23})^{\varepsilon}\left(Y^{(1)}_{123}\right)^{1/2-\varepsilon}I_{111}^{(1)}\,,$ (49)
making the coefficient in front of $J_{111}^{(1)}$ vanish in the differential equation for $\partial/\partial v_{11}J_{111}^{(1)}$ and also preserving the property that $J^{(1)}_{111}$ reduces to $J^{(0)}_{111}$ in the massless limit.
@@
$\frac{\partial J^{(0)}_{111}(\varepsilon-1)}{\partial v_{12}}=$ $(v_{12}v_{13}v_{23})^{\varepsilon}\left(Y^{(0)}_{123}\right)^{-1/2-\varepsilon}\frac{1-2\varepsilon}{2\,v_{12}}\Big{[}v_{12}(4-v_{12}-v_{13}-v_{23})I^{(0)}_{110}(\varepsilon-1)$
$+(2-v_{13})(v_{23}-v_{12}-v_{13})I^{(0)}_{101}(\varepsilon-1)+(2-v_{23})(v_{13}-v_{12}-v_{23})I^{(0)}_{011}(\varepsilon-1)\Big{]}.$ (50)
This constitutes a closed-form expression for ${\partial J^{(0)}_{111}(\varepsilon-1)}/{\partial v_{12}}$ since the two-denominator integrals are known analytically in terms of hypergeometric functions as well as in terms of an all-order $\varepsilon$-expansion.
For the derivative ${\partial J^{(1)}_{111}(\varepsilon-1)}/{\partial v_{11}}$ we find a similar closed-form expression. Here, it is convenient to express the right-hand side in terms of $I^{(1)}_{200}$ instead of $I^{(1)}_{100}$ to make spurious poles in $1/v_{11}$ cancel between coefficients, thus allowing for an integration starting at the massless point $v_{11}=0$. Using the identity
$I_{100}(v_{11};\varepsilon-1)=\frac{1}{2-2\varepsilon}\left[(3-2\varepsilon)I_{000}(\varepsilon-1)-v_{11}\,I_{200}(v_{11};\varepsilon-1)\right]$ (51)
to substitute $I_{100}$ for $I_{200}$, the derivative of $J_{111}^{(1)}$ becomes
$\frac{\partial J_{111}^{(1)}(\varepsilon-1)}{\partial v_{11}}=$ $2^{-1-\varepsilon}v_{23}^{\varepsilon}(2v_{12}v_{13}-v_{11}v_{23})^{-1+\varepsilon}\left(Y_{123}^{(1)}\right)^{-1/2-\varepsilon}$
×$\times\Big{[}(1-2\varepsilon)\Big{(}(v_{12}+v_{13}-v_{23})(v_{12}+v_{23}-v_{13}-v_{12}v_{23})+Y_{123}^{(1)}\Big{)}\,I_{110}^{(1)}(\varepsilon-1)$
$\quad+(1-2\varepsilon)\Big{(}(v_{12}+v_{13}-v_{23})(v_{13}+v_{23}-v_{12}-v_{13}v_{23})+Y_{123}^{(1)}\Big{)}\,I_{101}^{(1)}(\varepsilon-1)$
$\quad-(1-2\varepsilon)(v_{12}+v_{13}-v_{23})v_{23}(2-v_{23})\,I_{011}^{(0)}(\varepsilon-1)$
$\quad-2(1-v_{11})Y_{123}^{(1)}\,I_{200}^{(1)}(\varepsilon-1)-2(3-2\varepsilon)Y_{123}^{(1)}\,I_{000}^{(0)}(\varepsilon-1)\Big{]}.$ (52)
@@
VII Integration for master integrals
VII.1 Massless integral
To obtain the order $\varepsilon$ for $I_{111}^{(0)}$ in $d=4-2\varepsilon$ dimensions we only need the leading term of eq. (50) which is of order $\varepsilon^{0}$. Hence, we can set $\varepsilon=0$ in this equation and plug in the massless two-denominator integral in $d=6$ dimensions. Here we use the dimensionally-shifted massless two-denominator master integral given in eq. (106) of appendix E leading to
$\frac{\partial J^{(0)}_{111}(-1)}{\partial v_{12}}=\frac{\pi}{\sqrt{Y_{123}^{(0)}}}\left[\log\frac{v_{13}v_{23}}{4}-\log\frac{v_{12}}{2}+\frac{(v_{13}-v_{23})\log\frac{v_{13}}{v_{23}}}{v_{12}}-\frac{(2-v_{13}-v_{23})\log\frac{v_{12}}{2}}{2-v_{12}}\right]\equiv j^{(0)}(v_{12},v_{13},v_{23})\,.$ (53)
To obtain $J^{(0)}_{111}$ at a general kinematic point, we may choose the integration path $(1,1,1)\rightarrow(1,1,v_{23})\rightarrow(1,v_{13},v_{23})\rightarrow(v_{12},v_{13},v_{23})$ with boundary condition at the symmetric point $v_{12}=v_{13}=v_{23}=1$. The configuration of vectors used for the boundary value and the integration path are depicted in figure 2. By explicit calculation, the necessary boundary value is found to be
$J^{(0),d=6}_{111}(1,1,1)=I^{(0),d=6}_{111}(1,1,1)=2\pi\,\mathrm{C}\,,$ (54)

### v4

@@ 3 Mass reduction via two-point splitting
$\tilde{I}_{1,1,1}^{(3)}$ $(v_{12},v_{13},v_{23},v_{11},v_{22},v_{33})=\tilde{I}_{1,1,1}^{(1)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)},v_{33}\right)$
$-\tilde{I}_{1,1,1}^{(0)}\left(v_{(12)(13)},v_{(12)(31)},v_{(13)(31)}\right)+\tilde{I}_{1,1,1}^{(1)}\left(v_{1(12)},v_{1(13)},v_{(12)(13)},v_{11}\right)$
$+\tilde{I}_{1,1,1}^{(1)}\left(v_{3(12)},v_{3(31)},v_{(12)(31)},v_{33}\right)-\tilde{I}_{1,1,1}^{(0)}\left(v_{(21)(23)},v_{(21)(32)},v_{(23)(32)}\right)$
$+\tilde{I}_{1,1,1}^{(1)}\left(v_{2(21)},v_{2(23)},v_{(21)(23)},v_{22}\right)+\tilde{I}_{1,1,1}^{(1)}\left(v_{3(21)},v_{3(32)},v_{(21)(32)},v_{33}\right).$ (23)
4 Dimensional-shift identity
@@
Also the master integrals with $j_{i}=1$ and $v_{ii}=0$, which are collinearly divergent in $d=4$ dimensions, converge for sufficiently large dimension, i.e. small $\varepsilon$. In particular they are finite in $d=6$ dimensions. Hence applying eq.(24) to $I_{1,1,1}(\varepsilon)$ and subsequently reducing the integrals in six dimensions on the right to master integrals, results in all $1/\varepsilon$ poles being captured in prefactors of the master integrals.
As it turns out the identity connecting $I_{1,1,1}(\varepsilon)$ and $I_{1,1,1}(\varepsilon-1)$ simplifies further by shifting all integrals with two or less denominators back to $d=4-2\varepsilon$ dimensions. The necessary identities follow from eq.(24) by subsequently choosing zero, one, and two denominators, performing an IBP reduction and solving for the $d=6-2\varepsilon$ dimensional integrals. Going through the outlined steps, we find the remarkably simple relation
$I_{1,1,1}(\varepsilon)=\frac{1}{X_{123}}\left[X_{12\bar{3}}I_{1,1,0}(\varepsilon)+X_{1\bar{2}3}\,I_{1,0,1}(\varepsilon)+X_{\bar{1}23}\,I_{0,1,1}(\varepsilon)+\frac{2\varepsilon}{1-2\varepsilon}\,Y_{123}\,I_{1,1,1}(\varepsilon-1)\right].$ (25)
@@
Using the formula (25) has two benefits. Firstly for the purpose of analytical calculation, we gain the pole part and the $\varepsilon^{0}$ term “for free” since they are contained entirely in terms of known two-denominator integrals. Since $I_{1,1,1}(\varepsilon-1)$ is finite for $\varepsilon=0$ and is multiplied by an explicit $\varepsilon$ in eq. (25), it only starts contributing at order $\varepsilon$. Hence, for results to order $\varepsilon$ it suffices to calculate the leading term of $I_{1,1,1}(\varepsilon-1)$ by differential equations.
Secondly, from a numerical perspective, by expanding the integration measure of $I_{1,1,1}(\varepsilon-1)$ in a series in $\varepsilon$ we obtain a finite integral representation for every order in $\varepsilon$. In conjunction with the known expansion of the two-denominator integrals this can be used for numeric checks.
5 Differential operators for angular integrals
@@
Derivatives of the other invariants $v_{22}$, $v_{23}$, $v_{33}$ are trivially 0 due to the chosen ansatz. For $\partial/\partial v_{12}$, in the massless case, where $v_{11}=v_{22}=v_{33}=0$, we find
$\frac{\partial}{\partial v_{12}}=\frac{1}{Y_{123}^{(0)}}$ $\Big[(v_{13}+v_{23}-v_{12}-v_{13}v_{23})\vec{v}_{1}$
$-(2-v_{13})v_{13}\vec{v}_{2}+(v_{12}+v_{13}-v_{23}-v_{12}v_{13})\vec{v}_{3}\Big]\cdot\frac{\partial}{\partial\vec{v}_{1}},$ (39)
@@
$\frac{\partial}{\partial v_{11}}=\frac{1}{2Y_{123}^{(1)}}$ $\Big[-v_{23}(2-v_{23})\vec{v}_{1}+(v_{13}+v_{23}-v_{12}-v_{13}v_{23})\vec{v}_{2}$
$+(v_{12}+v_{23}-v_{13}-v_{12}v_{13})\vec{v}_{3}\Big]\cdot\frac{\partial}{\partial\vec{v}_{1}},$ (41)
@@
6 Differential equations
Having constructed suitable differential operators we are now equipped to derive differential equations for the three-denominator angular integral. The only master integral with genuinely three denominators is $I_{1,1,1}$. All other master integrals have at most two denominators and are known to all orders in $\varepsilon$. To construct the full solution for an arbitrary number of masses, we have to look at two cases. First we consider the massless integral. It is symmetric in its variables $v_{12},v_{13},v_{23}$, hence it suffices to only consider the differential equation $\partial/\partial v_{12}$ and construct a symmetric solution. As a boundary value we will use the integral at the symmetric point $v_{12}=v_{13}=v_{23}=1$. The second case to consider is the single-massive integral with $v_{11}\neq 0$. Here, we look at the differential equation with respect to $\partial/\partial v_{11}$ and use the massless integral as a boundary condition. As discussed in section 4 it is beneficial to consider the differential equations for integrals in $d=6-2\varepsilon$ dimensions. This way, the leading-order solution in $\varepsilon$ of the differential equation, which is of order $\varepsilon^{0}$, will provide a contribution to the order $\varepsilon$ of the master integral in $d=4-2\varepsilon$ dimensions.
@@ 6 Differential equations
$\frac{\partial I^{(0)}_{1,1,1}}{\partial v_{12}}=\frac{1}{Y_{123}^{(0)}}$ $\Big[v_{13}(v_{13}-v_{12}-v_{23})I^{(0)}_{2,1,1}+(v_{12}-v_{13}-v_{23}+v_{13}v_{23})I^{(0)}_{1,1,1}+v_{13}(2-v_{13})I^{(0)}_{2,0,1}$
$+(v_{23}-v_{12}-v_{13}+v_{12}v_{13})I^{(0)}_{2,1,0}\Big$] (43)
@@
$\frac{\partial I^{(1)}_{1,1,1}}{\partial v_{11}}=\frac{1}{2Y_{123}^{(1)}}$ $\Big[v_{23}(v_{23}-v_{12}-v_{13})I^{(1)}_{2,1,1}+v_{23}(2-v_{23})I^{(1)}_{1,1,1}+(v_{13}-v_{12}-v_{23}+v_{12}v_{23})I^{(1)}_{2,1,0}$
$+(v_{12}-v_{13}-v_{23}+v_{13}v_{23})I^{(1)}_{2,0,1}\Big$]. (44)
From here, we can use the recursion relations from appendix C to reduce the right-hand side of both equations to master integrals. In both cases, this results in a closed differential equation for $I_{1,1,1}$, where all other terms are expressed in terms of known lower-denominator integrals.
@@
$\frac{\partial I^{(0)}_{1,1,1}}{\partial v_{12}}$ $=\left[\frac{d-6}{2v_{12}}-\frac{d-5}{Y_{123}^{(0)}}\,(v_{13}+v_{23}-v_{23}-v_{13}v_{23})\right]I^{(0)}_{1,1,1}$
$+\frac{1}{2v_{12}Y_{123}^{(0)}}\Big[(d-5)v_{12}(4-v_{12}-v_{13}-v_{23})I^{(0)}_{1,1,0}$
$+(d-5)(2-v_{13})(v_{23}-v_{12}-v_{13})I^{(0)}_{1,0,1}+(d-5)(2-v_{23})(v_{13}-v_{12}-v_{23})I^{(0)}_{0,1,1}\Big$]. (45)
@@
$\frac{\partial I^{(1)}_{1,1,1}}{\partial v_{11}}$ $=\left[\frac{(6-d)v_{23}}{2(2v_{12}v_{13}-v_{11}v_{23})}+\frac{(d-5)v_{23}(2-v_{23})}{2Y_{123}^{(1)}}\right]I^{(1)}_{1,1,1}$
$+\frac{d-5}{2Y_{123}^{(1)}(2v_{12}v_{13}-v_{11}v_{23})}\Big[\Big((v_{12}+v_{13}-v_{23})(v_{12}+v_{23}-v_{13}-v_{12}v_{23})+Y_{123}^{(1)}\Big)I_{1,1,0}^{(1)}$
$+\Big((v_{12}+v_{13}-v_{23})(v_{13}+v_{23}-v_{12}-v_{13}v_{23})+Y_{123}^{(1)}\Big)I_{1,0,1}^{(1)}$
$-(v_{12}+v_{13}-v_{23})v_{23}(2-v_{23})I_{0,1,1}^{(0)}\Big]+\frac{3-d}{v_{11}(2v_{12}v_{13}-v_{11}v_{23})}I_{0,0,0}^{(0)}\,.$ (46)
We note that in both eqs. (45) and (46) all master integrals except for $I_{1,1,1}^{(0,1)}$ are known analytically in $d$ dimensions.
As discussed in section 4, it is most useful to consider the differential equations in $d=6-2\varepsilon$ dimensions. The standard procedure to approach differential equations of this form is bringing them to canonical $\varepsilon$-form, i.e. making the coefficient in front of $I_{1,1,1}$ proportional to $\varepsilon$. In the particular case at hand we can even make the entire coefficient disappear, leaving us with a right-hand side which is expressed solely in terms of known integrals This is achieved by rescaling the integral to $J_{1,1,1}=f(v)I_{1,1,1}$. Schematically writing the original eqs. (45) and (46) as
$\frac{\partial I_{1,1,1}}{\partial v}\,=\,c(v)\,I_{1,1,1}\,+\,r(v)\,$, (47)
@@
$\frac{\partial J_{1,1,1}}{\partial v}\,=\left(\frac{1}{f(v)}\frac{\partial f(v)}{\partial v}\,+\,c(v)\right)J_{1,1,1}\,+\,f(v)\,r(v)\,.$ (48)
@@
$J_{1,1,1}^{(0)}$ $\equiv(v_{12}v_{13}v_{23})^{\varepsilon}\left(Y^{(0)}_{123}\right)^{1/2-\varepsilon}I_{1,1,1}^{(0)}.$ (49)
The prefactor is chosen such that in the differential equation for $\partial/\partial v_{12}J_{1,1,1}^{(0)}$ the coefficient in front of $J_{1,1,1}^{(0)}$ becomes zero while simultaneously preserving the symmetry between variables. For the single-massive integral we rescale as
$J_{1,1,1}^{(1)}$ $\equiv\left(\frac{v_{23}}{2}\right)^{\varepsilon}(2v_{12}v_{13}-v_{11}v_{23})^{\varepsilon}\left(Y^{(1)}_{123}\right)^{1/2-\varepsilon}I_{1,1,1}^{(1)}\,$, (50)
making the coefficient in front of $J_{1,1,1}^{(1)}$ vanish in the differential equation for $\partial/\partial v_{11}J_{1,1,1}^{(1)}$ and also preserving the property that $J^{(1)}_{1,1,1}$ reduces to $J^{(0)}_{1,1,1}$ in the massless limit. We note that while it is clear that with the prefactor chosen this way $J^{(1)}_{1,1,1}$ reduces to $J^{(0)}_{1,1,1}$ before the $\varepsilon$-expansion, it also holds true in every order in the $\varepsilon$-expansion since $I_{1,1,1}^{(0)}$ is finite around six dimensions. All logarithmic divergences in the mass of the integral in $d=4-2\varepsilon$ are already explicit in the two-denominator contributions of eq. (25).
@@
$\frac{\partial J^{(0)}_{1,1,1}(\varepsilon-1)}{\partial v_{12}}=(v_{12}v_{13}v_{23})^{\varepsilon}\left(Y^{(0)}_{123}\right)^{-1/2-\varepsilon}\frac{1-2\varepsilon}{2\,v_{12}}\Big[v_{12}(4-v_{12}-v_{13}-v_{23})I^{(0)}_{1,1,0}(\varepsilon-1)$
$+(2-v_{13})(v_{23}-v_{12}-v_{13})I^{(0)}_{1,0,1}(\varepsilon-1)+(2-v_{23})(v_{13}-v_{12}-v_{23})I^{(0)}_{0,1,1}(\varepsilon-1)\Big].$ (51)
This constitutes a closed-form expression for ${\partial J^{(0)}_{1,1,1}(\varepsilon-1)}/{\partial v_{12}}$ since the two-denominator integrals are known analytically in terms of hypergeometric functions as well as in terms of an all-order $\varepsilon$-expansion.
For the derivative ${\partial J^{(1)}_{1,1,1}(\varepsilon-1)}/{\partial v_{11}}$ we find a similar closed-form expression. Here, it is convenient to express the right-hand side in terms of $I^{(1)}_{2,0,0}$ instead of $I^{(1)}_{1,0,0}$ to make spurious poles in $1/v_{11}$ cancel between coefficients, thus allowing for an integration starting at the massless point $v_{11}=0$. Using the identity
$I_{1,0,0}(v_{11};\varepsilon-1)=\frac{1}{2-2\varepsilon}\left[(3-2\varepsilon)I_{0,0,0}(\varepsilon-1)-v_{11}\,I_{2,0,0}(v_{11};\varepsilon-1)\right]$ (52)
to substitute $I_{1,0,0}$ for $I_{2,0,0}$, the derivative of $J_{1,1,1}^{(1)}$ becomes
$\frac{\partial J_{1,1,1}^{(1)}(\varepsilon-1)}{\partial v_{11}}=$ $2^{-1-\varepsilon}v_{23}^{\varepsilon}(2v_{12}v_{13}-v_{11}v_{23})^{-1+\varepsilon}\left(Y_{123}^{(1)}\right)^{-1/2-\varepsilon}$
$\times\Big[(1-2\varepsilon)\Big((v_{12}+v_{13}-v_{23})(v_{12}+v_{23}-v_{13}-v_{12}v_{23})+Y_{123}^{(1)}\Big)\,I_{1,1,0}^{(1)}(\varepsilon-1)$
$\quad+(1-2\varepsilon)\Big((v_{12}+v_{13}-v_{23})(v_{13}+v_{23}-v_{12}-v_{13}v_{23})+Y_{123}^{(1)}\Big)\,I_{1,0,1}^{(1)}(\varepsilon-1)$
$\quad-(1-2\varepsilon)(v_{12}+v_{13}-v_{23})v_{23}(2-v_{23})\,I_{0,1,1}^{(0)}(\varepsilon-1)$
$\quad-2(1-v_{11})Y_{123}^{(1)}\,I_{2,0,0}^{(1)}(\varepsilon-1)-2(3-2\varepsilon)Y_{123}^{(1)}\,I_{0,0,0}^{(0)}(\varepsilon-1)\Big].$ (53)
@@
7 Integration for master integrals
7.1 Massless integral
To obtain the order $\varepsilon$ for $I_{1,1,1}^{(0)}$ in $d=4-2\varepsilon$ dimensions we only need the leading term of eq. (51) which is of order $\varepsilon^{0}$. Hence, we can set $\varepsilon=0$ in this equation and plug in the massless two-denominator integral in $d=6$ dimensions. Here we use the dimensionally-shifted massless two-denominator master integral given in eq. (107) of appendix E leading to
$\frac{\partial J^{(0)}_{1,1,1}(-1)}{\partial v_{12}}$ $=\frac{\pi}{\sqrt{Y_{123}^{(0)}}}\left[\log\frac{v_{13}v_{23}}{4}-\log\frac{v_{12}}{2}+\frac{(v_{13}-v_{23})\log\frac{v_{13}}{v_{23}}}{v_{12}}-\frac{(2-v_{13}-v_{23})\log\frac{v_{12}}{2}}{2-v_{12}}\right]$
$\equiv j^{(0)}(v_{12},v_{13},v_{23})\,.$ (54)
To obtain $J^{(0)}_{1,1,1}$ at a general kinematic point, we may choose the integration path $(1,1,1)\rightarrow(1,1,v_{23})\rightarrow(1,v_{13},v_{23})\rightarrow(v_{12},v_{13},v_{23})$ with boundary condition at the symmetric point $v_{12}=v_{13}=v_{23}=1$. The configuration of vectors used for the boundary value and the integration path are depicted in figure 2. By explicit calculation, the necessary boundary value is found to be
$J^{(0),d=6}_{1,1,1}(1,1,1)=I^{(0),d=6}_{1,1,1}(1,1,1)=2\pi\,\mathrm{C}\,$, (55)